OpenStudy (anonymous):

For what values of x does the graph of f have a horizontal tangent? (Use n as your integer variable. Enter your answers as a comma-separated list.) f(x) = x − 2 sin x

OpenStudy (anonymous):

solve f'(x) = 0

OpenStudy (anonymous):

f'(x) + cos(x) = 0 ?

OpenStudy (anonymous):

I have ten minutes could someone give me a clear answer?

OpenStudy (anonymous):

f'(x) = 1 - 2 cos x

OpenStudy (anonymous):

1 - 2 cos x = 0 cos x = 1/2

OpenStudy (anonymous):

is the answer 0, 1/2 then?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

use unit circle http://www.sciencedigest.org/UnitCircle.gif

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

the answer is Pi/3, 5*Pi/3

OpenStudy (anonymous):

+ all multiples of 2Pi

OpenStudy (anonymous):

cos x = 1/2 means you looks for the angle where the x coordinate is 1/2

OpenStudy (anonymous):

so like pi/6?

OpenStudy (anonymous):

no that's a solution to sin x = 1/2

OpenStudy (anonymous):

I'm really lost. I really need to sleep more. 4 mins to go, feeling generous and just want to throw the full answer out there?

OpenStudy (anonymous):

oh, cos is on the left side, sin on the right side. Like i said i need more sleep. heh.

OpenStudy (anonymous):

Pi/3 + n*2*Pi and 5*Pi/3 + n*2*Pi, not sure how they want you to enter it

OpenStudy (anonymous):

well, I ran out of time, but I appreciate the effort anyway thanks! here's your medal